# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 2.5. Graphical Representation: Xent Games as Diagrams
In order to study simply Xent Games, especially the ones written in XGL, it is often convenient to express their operational logic via graphical diagrammatic representations: this is the way that the examples of Section 3 are represented.
@@
  • It is also a good way to grasp the nature of the Xent Game space, in particular the interactions between the string space operations, the xent operations and the player’s moves is much easier to understand.
  • While the assembly-like nature of XGL is usually easy to follow, it is often desirable to eliminate some redundancies involved with the notation (swapping the name of two registers leaves a game unchanged, it is not clear which registers are in fact used), which the graphical representation does.
The philosophy of XGD is quite simple, and in some sense closer to the theoretical move description (see Section 2.2.5). The XGD representation is based on symbols linked via edges of various stroke styles, surrounded by constraint boxes. See Figure 2.1 for a reference of the symbols, and Figures 3.2, 3.4, 3.6, 3.8, and 3.11 for some simple examples of Xent Games (note that it is easier to understand XGD by looking at these examples first).
@@ 2.5.1. Colors and String Space Variables
  • Associated with each player is a color and a negative color: the players $black$, $white$, $alice$, $bob$, $carol$, $david$, and env have the (positive) colors $black$, $white$, $red$, $green$, $blue$ $and$ purple respectively, and they have the negative colors $white$, $black$, $cyan$, $magenta$, $yellow$, and $olive$ respectively.
@@
    • A colored circle $\circ$ for $elicit$ inputs, with a number inside the circle
@@ 2.5.2. Xent Symbols
  • For the xent symbols that are not in an $ensure$ box (see below), for instance:
@@
  • Each xent symbol in an $ensure$ box is colored similarly, and gives a constraint on the move of the $elicit$ contained in the same box: the sum of the xent symbols in the box in favor of the player elicited should be positive.
@@
  • Using the relevant colors and negative colors (see Figure 2.1).
@@ 2.5.3. Elicit and Reveals
As discussed above the $elicit$ circle boxes are represented by circles.
@@
  • Inside of the elicit circles, numbers can be written which specify the number of tokens being elicited.
  • At the bottom right of an elicit, a number can appear, denoting the maximal length of the elicit.
  • The revealed information if we elicit a red/green/blue player is represented by star of the relevant color appearing to the left (i.e. having been seen) of the $elicit$ box.
@@ 2.5.4. Ensure Boxes
The $ensure$ constraints are represented by referring to elicit symbols. An $elicit$ can have several corresponding ensure statements; in case an $ensure$ contains several $elicit$ statements, the one to which one comes back if the $ensure$ fails is surrounded by a circle.
@@
The $ensure$ boxes are implicitly linked by strokes to at least an elicit box, and if they fall back on the last elicit if the $ensure$ fails.
@@ 2.5.5. Cat/Cut
  • The concatenation of two strings (with a default spacing character between them) is represented by an edge linking the two corresponding nodes oriented (with an arrow symbol) according to which string appears before which one ($s\to t$ represents $s+t$, i.e. $s$ followed by $t$).
  • Each cat/cut operation corresponds to a certain stroke type (e.g. regular, dashed, wiggly, doubled, …)
@@
  • The cut of a string $s$ by another $t$ is represented by a single-arrow oriented edge being ‘intercepted’ by a double-arrow oriented edge (with same stroke type): the emerging double-arrow edge carries $s//t$, while the emerging single-arrow edge carries $s\%t$.
  • A string can belong to e.g. several concatenations with several edges being incident if they are represented using different stroke styles.
  • An unoriented edge with a certain stroke type ‘carries’ the result of the string operation ‘performed’ by the corresponding stroke type to an operation (e.g. a xent symbol).
  • A string can be stored in a star if there is an unoriented edge ‘carrying’ the string towards a star; there should be exactly one unoriented edge incident to a star.
  • Dotted (unoriented) edges carry values from a node to another, with the oriented being from left to right.
@@ 3. Xent Game Examples
  • In Section 3.2, a number of two-player perfect information zero-sum games are presented: informally, these are minimax games, i.e. since each player’s gain is the other player’s loss, theoretically the optimal play is to minimize the other player’s return over all their possible actions following one’s action.
@@ 3.1. One-Player Combinatorial Games
In Figure 3.1, we present three human-friendly optimization games, which involve the $xed$ function (with $xed\left(s|t\right)=xent\left(s\right)-xent\left(s|t\right)$). Note that the first game is plausibly something that would be a desirable output from an LLM, but that it is actually hard to solve (taking a small temperature leads to a greedy approximation of it). Through substantial effort to try on human players,
@@
In Figure 3.3, we present a small collection of games each of which comes with slightly different interpretation, despite fairly short codes and limited number of involved variables.
@@ 3.2.2. Chess
As described in Figure 3.7, a naive way to play chess using Xent Games is to play 20 moves and then as a judge model to decide who is winning, and use the cross-entropy difference between the estimates as a way; the judge must enforce the validity of moves at every step. Alternatively, the full game of chess can be implemented, but since the theoretical upper bound on chess game lengths is around 5000, the game must run for all these steps.
@@ 3.2.3. Naive Sprig: Concise Mathematical Proof Debate
assign(s=story("Get a provable elementary mathematical statement"),s3=story("Get three typical one-line math arguments"))
@@
assign(s0=s0+"white proposes to use the following prove the last statement:" + t)
@@
reward(white, xent(t0 + "is a correct elementary mathematical proof w/o missing details: false"))
reward(black, xent(t0 + "is a correct elementary mathematical proof w/o missing details: true"))
@@ 3.3. Imperfect Information Games
Beyond the perfect information setup, the imperfect information games bring considerably more sophistication for the same space of allowed moves. We present a few games that are illustrative of the challenges associated with imperfect information, involving in particular information sharding, coordination, guessing, and bluff.
@@ 3.3.1. Secret Sharing
A simple example of an imperfect-information game is secret sharing (Figure 3.10), where Alice must split a secret into Carol’s and David’s share so that neither can find the secret, but so that Bob who has both shares can find it. Notice that this kind of problems have simple cryptographic solutions, but that playing such a game with LLMs can give interesting results.
@@ 3.3.2. Coordination
Another example of imperfect information game is coordination (Figure 3.12), where Bob and Carol have access to the same information, and must produce the closest possible guesses.
@@ 4.1. Xent Game Eval Mode
By definition a Xent Game $G$, beyond its game operation code, involves some metadata $D$ including ‘linking’ metadata:
@@
When using a Xent Game to evaluate a model $\mathcal{M}$, the idea is to ‘freeze’ a game $\left(G,D\right)$, pick an active player $P\in\mathcal{A}\left(G\right)$ in the game $G$ and to consider the family of games $\left(G,D|_{P=\mathcal{M}}\right)$ where $D|_{P=\mathcal{M}}$ denotes the linking of $P$ with the model $\mathcal{M}$, overriding the linking of $\mathcal{P}$ in $D$, i.e. forcing the player $P$ to become the main character played by $\mathcal{M}$. Informally, this corresponds to the following simple idea: consider a game $G$, pick a player $P$ and measure how the model plays the role of the ‘character’ $P$.
Typically, running a Xent Game in Eval Mode furthermore involves specifying a random ‘map-seed’ $\sigma$ used to set the instance random data (unless we deal with a game with deterministic instances, e.g. like Chess or Go), and in the outputs of the players except the model being evaluated; we typically use a different ‘play-seed’ $\varsigma$.
@@
In order to allow for unified normalization across various games, game maps and game settings, the game scores should be normalized using a base model (see Section 5.2 below).
@@ 4.1.2. Well-Posedness
A notion that is related to the one discussed in Section 4.1.1 above is that of well-posedness. This can be viewed as an analog to the notion of boundedness in optimization: trying to maximize a function that is not bounded from above cannot yield meaningful results, even if in practice values may be bounded by machine precision (whatever result we will get will reflect something about the machine running the computations rather than about the problem). In the setting of Xent Games, rewards on any game instance are naturally bounded by the bounded sizes of the strings, but that does not imply solutions are anyhow meaningful, as they may be more reflective of irrelevant model artifacts rather than anything else.
@@
For $\epsilon>0$, we say a Xent Game $G$ (in eval mode) is $\epsilon$-well-posed if the main player’s optimal strategy is not affected by $\epsilon$-clipping of probabilities, i.e. replacing $x\mapsto-\log x$ by $x\mapsto-\max\left(\log x,\log\epsilon\right)$ in the cross-entropy.
@@ 4.1.3. Game Repetition
An important feature of learning to play games is repetition, i.e. learning from mistakes, and more generally understanding the interplay between a game play history and the rules of the games to optimize one’s strategy, by putting all of this information as a context for the model to play. A useful setting in practice is to allow the to replay of games by the main character played by the model $\mathcal{M}$, while keeping the same map-seed and the same play-seed over the various replays, giving as context the past games to $\mathcal{M}$, while keeping no past game context to the other NPC agents. This setup, which advantages $\mathcal{M}$ over the NPC players, can be expected to give an increased score as the attempts are repeated for a reasonable class of games: as discussed in Section 4.2.1 below, we call these the few-shot playable games.
@@ 4.1.4. Game Fine-Tuning
While we intentionally avoid setting specific details about the process, it is useful to take as an example the one developed in Section 4.3.2. Note that we do not require that the model has actually gotten better at $G$: as will be discussed in Section 4.2 below, this corresponds to the family of fine-tuning playable games.
It should be noted that a fine-tuning mechanism should not be assumed to exist for all models $\mathcal{M}$: for instance, some closed models can be just available in inference mode and not be fine-tunable. In this paper, we simply required the existence of a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$ for which a fine-tuning mechanism is available, that makes enough games playable (see Section 6 below).
@@ 4.2. Playability: Few-Shot and Fine-Tuning Definitions
Central to our investigation of well-posed Xent Games is the notion of playability, which informally corresponds to a quality of certain games that models can get improved performance on, either by repeated play context (few-shot definition) or by reinforcement learning (fine-tuning definition). This should be thought of in opposition to non-playable games, which contain for instance as a non-trivial example the breaking of a secure code (even with a small secret length), as one either succeeds or fails and there is no intermediate reward and no sense that we get closer to succeeding.
@@ 4.2.1. Few-Shot Definition
A remarkable ability of modern LLMs is their so-called few-shot learning abilities [OAI2020, KGRMI22]: from a limited number of feedbacks directly put in their context window, models are often able to substantially improve their responses. The notion of few-shot playability by a model $\mathcal{M}$ refers to those games where the few-shot abilities are reflected in game-play:
@@
For a game $G$ and $N$ random samples (of map seeds and play seeds), we define the average running max score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ as the (random) sequence defined by
@@
A well-posed game $G$ is $\left(N,\epsilon,m\right)$-playable in few-shot mode by a model $\mathcal{M}$, if with probability $1$-$\epsilon$, the score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ is strictly increasing for at least $m$ steps. We say that a playable game is $\left(N,\epsilon,\theta,m\right)$ playable if there exists $m$ thresholds $k_{1}<\cdots<k_{m}$ such that $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ exceeds $b+j\theta$ for the first time at $k_{j}$ for $j=1,\ldots,m$ where $b$ is the expected score at the first iteration.
@@
This definition means that in practice, averaging over a sufficiently large number of samples, we do see the running max score increase substantially for a while (i.e. at every step, the model has done better in one sample). We choose this definition rather than that of an expectation as the latter is sensitive to very unlikely (e.g. breaking a secure cipher is not a playable game, yet the probability of cracking it is technically nonzero at every step).
As we will see, in Section 5, few-shot playability is an important a priori property for benchmarking: the benchmarking score of a game will be defined relative to the times when a base model $\mathcal{M}_{\mathcal{B}}$ crosses $\theta$-levels for a game’s repeated play.
@@ 4.2.2. Numerical Example: Benchmarking with Single-Text
The initial benchmarking results have a number of interesting features. That being said, the results clearly demonstrate that current models are able to improve their performance on a well-posed game when given the chance to provide new answers after seeing the results of their previous choices.
The following plots show performance of GPT-4.1, Claude Opus 4, Gemini 2.5 Pro, and Grok 3 when playing the single-text game posed in the Xent Game Language in an iterated scenario using GPT2 as the judge model. Each test case is the single text game with a different generated string. While these specific test cases show clearly improvement across the set of test models, often results would be more mixed, showing flat or even occasionally negative performance given more information.
Figure 4.1. Performances for four flagship models playing single text Xent Game (with $N=20$, running for $32$ iterations): the top plot shows the averaged scores (see Expression 4.1) and the bottom plot shows the $arms$ (see Definition 33). We see that the $arms$ keeps improving for all models up to $k=17$.

### v2

@@ 2.5. Graphical Representation: Xent Games as Diagrams
In order to study simple Xent Games, especially the ones written in XGL, it is often convenient to express their operational logic via graphical diagrammatic representations: this is the way that the examples of Section 3 are represented.
@@
  • It is also a good way to grasp the nature of the Xent Game space, in particular the interactions between the string space operations, the xent operations, and the player’s moves are much easier to understand.
  • While the assembly-like nature of XGL is usually easy to follow, the graphical representation eliminate some redundancies involved with the notation (e.g. swapping the name of two registers leaves a game unchanged, it is not clear which registers are in fact used).
The philosophy of XGD is quite simple and in some sense closer to the theoretical move description (see Section 2.2.5). The XGD representation is based on symbols linked via edges of various stroke styles, surrounded by constraint boxes. See Figure 2.1 for a reference of the symbols, and Figures 3.2, 3.4, 3.6, 3.8, and 3.11 for some simple examples of Xent Games. Note that it is easier to understand XGD by looking at these examples first.
@@ 2.5.1. Colors and String Space Variables
  • Associated with each player is a color and a negative color: the players $black$, $white$, $alice$, $bob$, $carol$, $david$, and env have the (positive) colors $black$, $white$, $red$, $green$, $blue$, $and$ purple respectively, and they have the negative colors $white$, $black$, $cyan$, $magenta$, $yellow$, and $olive$ respectively.
@@
    • A colored circle $\circ$ for $elicit$ inputs, with a number inside the circle representing max length
@@ 2.5.2. Xent Symbols
  • For the xent symbols that are not in an $ensure$ box (see below):
@@
  • Each xent symbol in an $ensure$ box is colored similarly, and gives a constraint on the move of the $elicit$ contained in the same box: the sum of the xent symbols in the box must be positive for the $ensure$ to pass.
@@
  • Each xent symbol is associated with a player via their color and negative color (see Figure 2.1).
@@ 2.5.3. Elicit and Reveals
As discussed above $elicit$ instructions are represented by circles.
@@
  • Inside of the elicit circles, numbers can be written, denoting the maximal length of the elicit.
  • The revealed information, if we elicit a non-omniscient player, is represented by star of the relevant color appearing to the left (i.e. having been seen) of the $elicit$ box.
@@ 2.5.4. Ensure Boxes
The $ensure$ constraints are represented by referring to elicit symbols. An $elicit$ can have several corresponding ensure statements; in the case an $ensure$ contains several $elicit$ statements, the $elicit$ statement which execution jumps back to if the $ensure$ fails is surrounded by a circle.
@@
The $ensure$ boxes are implicitly linked by strokes to at least one $elicit$ box, and if they fall back on the last elicit if the $ensure$ fails.
@@ 2.5.5. Cat/Cut
  • The concatenation of two strings (with a default spacing character between them) is represented by an edge linking the two corresponding nodes oriented with an arrow symbol, determining the order of the concatenation ($s\to t$ represents $s+t$, i.e. $s$ followed by $t$).
  • Each cat/cut operation corresponds to a certain stroke type (e.g. regular, dashed, wiggly, doubled, etc.)
@@
  • The cut of a string $s$ by another $t$ is represented by a single-arrow oriented edge being ‘intercepted’ by a double-arrow oriented edge (with same stroke type). The emerging double-arrow edge carries $s//t$, while the emerging single-arrow edge carries $s\%t$.
  • A string can belong to several concatenations with several edges being incident if they are represented using different stroke styles.
  • An unoriented edge with a certain stroke type ‘carries’ the result of the string operation ‘performed’ by the corresponding stroke type to another operation (e.g. a xent symbol).
  • A string can be stored in a star if there is an unoriented edge ‘carrying’ the string towards the star; there should be exactly one unoriented edge incident to a star.
  • Dotted (unoriented) edges carry values from a node to another, with the orientation being from left to right.
@@ 3. Xent Game Examples
  • In Section 3.2, a number of two-player perfect information zero-sum games are presented: informally, these are minimax games, i.e. since each player’s gain is the other player’s loss, theoretically the optimal play is to minimize the other player’s return over all their possible actions following one’s own action.
@@ 3.1. One-Player Combinatorial Games
In Figure 3.1, we present three human-friendly optimization games which involve the $xed$ function (with $xed\left(s|t\right)=xent\left(s\right)-xent\left(s|t\right)$). Note that the first game leads to results that would plausibly be a desirable output from an LLM while being quite hard to solve in practice (taking a small temperature leads to a greedy approximation of the solution).
@@
In Figure 3.3, we present a small collection of games each of which comes with a slightly different interpretation, despite fairly short codes and limited number of involved variables.
@@ 3.2.2. Chess
As described in Figure 3.7, a naive way to play chess using Xent Games is to play 20 moves and and use the judge model’s cross-entropy difference between the estimates for ‘white is winning’ vs ‘black is winning’ to determine the winner. The judge must enforce the validity of moves at every step. Alternatively, the full game of chess can be implemented, but since the theoretical upper bound on chess game lengths is around 5000, the game must be allowed to run for all these steps.
@@ 3.2.3. Naive Sprig: Concise Mathematical Proof Debate
assign(s=story("Get a provable elementary mathematical statement"))
assign(s3=story("Get three typical one-line math arguments"))
@@
assign(s0=s0+"white proposes the following prove the last statement:" + t)
@@
reward(white, xent(t0 + "is a correct math proof w/o missing details: false"))
reward(black, xent(t0 + "is a correct math proof w/o missing details: true"))
@@ 3.3. Imperfect Information Games
Beyond the perfect information setup, the imperfect information games bring considerably more sophistication for the same space of allowed moves. We present a few games that are illustrative of the challenges associated with imperfect information, involving in particular information sharing, coordination, guessing, and bluff.
@@ 3.3.1. Secret Sharing
A simple example of an imperfect-information game is secret sharing (Figure 3.10), where Alice must split a secret into Carol’s and David’s share so that neither can find the secret, but so that Bob, who has both shares, can find it. Notice that these kind of problems have simple cryptographic solutions, but that playing such a game with LLMs can still give interesting results.
@@ 3.3.2. Coordination
Another example of an imperfect information game is coordination (Figure 3.12), where Bob and Carol have access to the same information and must produce the closest possible guesses.
@@ 4.1. Xent Game Eval Mode
By definition, a Xent Game $G$, beyond its game operation code, involves some metadata $D$ including ‘linking’ metadata:
@@
When using a Xent Game to evaluate a model $\mathcal{M}$, the idea is to ‘freeze’ a game $\left(G,D\right)$, pick an active player $P\in\mathcal{A}\left(G\right)$ in the game $G$ and to consider the family of games $\left(G,D|_{P=\mathcal{M}}\right)$ where $D|_{P=\mathcal{M}}$ denotes the linking of $P$ with the model $\mathcal{M}$, overriding the linking of $\mathcal{P}$ in $D$, i.e. forcing the player $P$ to become the main character played by $\mathcal{M}$. Informally, this corresponds to the following simple idea: consider a game $G$, pick a player $P$, and measure how the model plays the role of the ‘character’ $P$.
Running a Xent Game in Eval Mode furthermore involves specifying a random ‘map-seed’ $\sigma$ used to set the instance random data (unless we deal with a game with deterministic instances, e.g. Chess or Go), and in the generation of the outputs of the players except the model being evaluated; for the model being evaluated, we typically use a different ‘play-seed’ $\varsigma$.
@@
In order to allow for unified normalization across various games, game maps, and game settings, the game scores should be normalized using a base model (see Section 5.2 below).
@@ 4.1.2. Well-Posedness
A notion that is related to the one discussed in Section 4.1.1 above is that of well-posedness. This can be viewed as an analog to the notion of boundedness in optimization: trying to maximize a function that is not bounded from above cannot yield meaningful results, even if, in practice, values may be bounded by machine precision (whatever result we will get will reflect something about the machine running the computations rather than about the problem). In the setting of Xent Games, rewards on any game instance are naturally bounded by the bounded sizes of the strings, but that does not imply solutions are anyhow meaningful, as they may be more reflective of irrelevant model artifacts rather than anything else.
@@
For $\epsilon>0$, we say a Xent Game $G$ (in eval mode) is $\epsilon$-well-posed if the main player’s optimal strategy is not affected by $\epsilon$-clipping of probabilities, i.e. replacing $x\mapsto-\log x$ by $x\mapsto-\max\left(\log x,\log\epsilon\right)$ in the cross-entropy computation.
@@ 4.1.3. Game Repetition
An important feature of learning to play games is repetition, i.e. learning from mistakes, and more generally understanding the interplay between a game play history and the rules of the games to optimize one’s strategy. A useful setting in practice is to allow the to replay of games by the main character played by the model $\mathcal{M}$, while keeping the same map-seed and the same play-seed over the various replays, giving the past games as context to $\mathcal{M}$, while keeping no past game context to the other NPC agents. This setup, which advantages $\mathcal{M}$ over the NPC players, can be expected to give an increased score as the attempts are repeated for a reasonable class of games: as discussed in Section 4.2.1 below, we call these the few-shot playable games.
@@ 4.1.4. Game Fine-Tuning
While we intentionally avoid setting specific details about the process, it is useful to take as an example the one developed in Section 4.3.2 below. Note that we do not require that the model has actually gotten better at $G$: as will be discussed in Section 4.2 below, this corresponds to the family of fine-tuning playable games.
It should be noted that a fine-tuning mechanism should not be assumed to exist for all models $\mathcal{M}$: for instance, some closed models are only available in inference mode and not be fine-tunable. In this paper, we simply required the existence of a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$ for which a fine-tuning mechanism is available that makes enough games playable (see Section 6 below).
@@ 4.2. Playability: Few-Shot and Fine-Tuning Definitions
Central to our investigation of well-posed Xent Games is the notion of playability, which informally corresponds to a quality of certain games that models can get improved performance on, either by repeated play context (few-shot definition) or by reinforcement learning (fine-tuning definition). This should be thought of in opposition to non-playable games, which contain, for instance, as a non-trivial example, the breaking of a secure code (even with a small secret length), as one either succeeds or fails and there is no intermediate reward and no sense that we get closer to succeeding.
@@ 4.2.1. Few-Shot Definition
A remarkable ability of modern LLMs is their so-called few-shot learning abilities [OAI2020, KGRMI22]: from a limited number of feedback examples directly put in their context window, models are often able to substantially improve their responses. The notion of few-shot playability by a model $\mathcal{M}$ refers to those games where the few-shot abilities are reflected in game-play:
@@
For a game $G$ and $N$ random samples of map seeds and play seeds, we define the average running max score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ as the (random) sequence defined by
@@
A well-posed game $G$ is $\left(N,\epsilon,m\right)$-playable in few-shot mode by a model $\mathcal{M}$ if, with probability $1$-$\epsilon$, the score $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ is strictly increasing for at least $m$ steps. We say that a playable game is $\left(N,\epsilon,\theta,m\right)$ playable if there exists $m$ thresholds $k_{1}<\cdots<k_{m}$ such that $\mathrm{arms}_{N,k}\left(G,\mathcal{M}\right)$ exceeds $b+j\theta$ for the first time at $k_{j}$ for $j=1,\ldots,m$ where $b$ is the expected score at the first iteration.
@@
This definition means that, in practice, averaging over a sufficiently large number of samples, we do see the running max score increase substantially for a while (i.e. at every step, the model has done better in one sample). We choose this definition rather than that of an expectation as the latter is sensitive to very unlikely outcomes (e.g. breaking a secure cipher is not a playable game, yet the probability of cracking it is technically nonzero at every step).
As we will see in Section 5, few-shot playability is an important a priori property for benchmarking: the benchmarking score of a game will be defined relative to the times when a base model $\mathcal{M}_{\mathcal{B}}$ crosses $\theta$-levels for a game’s repeated play.
@@ 4.2.2. Numerical Example: Benchmarking with Single-Text
The plots in Figure 4.1 show the performances of GPT-4.1, Claude Opus 4, Gemini 2.5 Pro, and Grok 3 when playing the single-text game posed in the Xent Game Language in an iterated scenario using GPT2 as the judge model. Each test case is the single text game with a different generated string. While these specific test cases show clear improvement across the set of test models, individual game map results are more mixed, showing flat or even occasionally negative performance as iterations increased. Our results are obtained from the code available in the xega repository on GitHub.
That being said, the results clearly demonstrate that current models are able to improve their performance on a well-posed game when given the chance to provide new answers after seeing the results of their previous choices.
Figure 4.1. Performances for four flagship models playing the single text Xent Game (with $N=20$, running for $32$ iterations): the top plot shows the averaged scores (see Expression 4.1) and the bottom plot shows the $arms$ (see Definition 33). We see that the $arms$ keeps improving for all models up to $k=17$.